Solve the system by substitution
x + 7y = -2
3x + y =34

Answer by privani(34)

x + 7y = -2
x = -2 - 7y .............(1)
3x + y =34 .............(2)
Substitute (1) in (2)
3x + y =34
3(-2-7y)+y=34
-6-21y+y=34
-20y=34+6
-20y=40
y=40/-20
y=-2
Substitute y=-2 in (1)
x=-2-7y
x=-2-7(-2)
x=-2+14
x=12